Flying against the wind, an airplane travels 5390 kilometers in 7 hours. Flying with the wind, the same plane travels 3750 kilometers in 3 hours. What is the rate of the plane in still air and what is the rate of the wind?
Answer by ikleyn(52835)

.
Flying against the wind, an airplane travels 5390 kilometers in 7 hours.
Flying with the wind, the same plane travels 3750 kilometers in 3 hours.
What is the rate of the plane in still air and what is the rate of the wind?
~~~~~~~~~~~~~~~

Let  u  be the airplane rate in still air (in miles per hour), and
let  v  be the rate of the wind.


Then the airplane effective speed flying with the wind is  u+v  miles per hour,
while its effective speed flying against      the wind is  u-v  miles per hour.


From the condition, we have this system of equstions


    u + v = 3750/3 = 1250  mph    (1)

    u - v = 5390/7 =  770  mph    (2)


To solve equations, add them. You will get


    2u = 1250 + 770 = 2020;  hence  u = 2020/2 = 1010  mph.


Then from equation (1),

    v = 1250 - u = 1250 - 1010 = 240 mph.


ANSWER.  The rate of the airplane in still air is 1010 miles per hour;  the rate of the wind is 240 mph.

Solved.


/////////////

Post-solution note


        The rate of the wind seems TOOOOO high to be realistic.

        It is twice the air speed in the eye of an hurricane.

        So keep in mind please that your input produces unrealistic output.

-------------

It is a typical "tailwind and headwind" word problem.

See the lessons
    - Wind and Current problems
    - Wind and Current problems solvable by quadratic equations
    - Selected problems from the archive on a plane flying with and against the wind
in this site, where you will find other similar solved problems with detailed explanations.


Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook under the section "Word problems",  the topic "Travel and Distance problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.